# Time-resolved diamond magnetic microscopy of superparamagnetic iron-oxide nanoparticles

### v1

@@ Introduction
Here, we use diamond magnetic microscopy to measure the field-dependent magnetization and time-domain Néel relaxation of over a hundred individual SPIONs in parallel. We compile histograms and extract statistics of key SPION parameters, such as saturation magnetic moment and Néel relaxation time, to characterize the ensemble distribution. The data reveal remarkable properties, including starkly steeper magnetization curves for some SPIONs compared to those obtained from ensemble methods. Such a rich heterogeneity is masked in ensemble methods, suggesting our method can produce complementary insights in fundamental studies of nanomagnetism.
@@
For diamond magnetic microscopy of SPIONs, we used the experimental setup depicted in Fig. 1(a) (see also Appendix I). The diamond is a $100\mbox{-}{\rm\upmu m}$-thick electronic-grade membrane with faces polished normal to a [110] crystal direction such that two of the NV axes lie in plane. Ion implantation and annealing is used to form a ${\sim}150~{}{\rm nm}$-thick NV layer near the diamond surface on which SPION samples are dispersed (Appendix II). A beam of 532-nm light continuously illuminates NV centers in a $45\times 45~{}{\rm\upmu m}^{2}$ field of view (FOV) via an oil-immersion objective. NV center fluorescence is collected through the same objective, spectrally filtered (passing $663{\mbox{-}}800~{}{\rm nm}$), and imaged onto a scientific CMOS image sensor, with a lateral resolution of ${\sim}350~{}{\rm nm}$. The diamond is mounted on a glass coverslip using UV-curing adhesive. Microwaves are delivered to the NV centers via copper traces on the coverslip. The diamond is positioned so that one of the in-plane NV axes lies parallel to an applied magnetic field $\vec{B_{0}}$, generated by a pair of coils. The diamond-coverslip assembly is mounted on a 3-axis piezo-actuated translation stage. Fluorescence images of fiducial markers milled into the diamond NV layer are used to adjust the translation stage and stabilize the diamond position [36], limiting drift to ${\lesssim}100~{}{\rm nm/day}$.
@@
The SPIONs (Ocean Nanotech SOR30) have Fe3O4 magnetic cores and an oleic acid coating, with a manufacturer-specified diameter of ${\sim}30~{}{\rm nm}$. SPION suspensions are diluted in hexane, drop-cast onto the diamond, and dried in air, resulting in isolated single SPIONs with an average spacing of ${\sim}3~{}{\rm\upmu m}$ (Appendix III.1). SPIONs lying within the diamond magnetic microscope’s FOV are identified by scanning electron microscopy (SEM) and co-localized with respect to the fiducial markers. In the specific FOV presented here (Appendix III), we identified and localized $259$ single SPIONs, of which $101$ were sufficiently isolated (${\gtrsim}1.5~{}{\rm\upmu m}$ from nearest neighbor and image edges) such that their magnetic dipole patterns could be individually resolved with negligible overlap in our diamond magnetic microscope. Figure 1(b) shows a SEM image of five isolated SPIONs on the diamond surface within a sub-region of this FOV.
@@
Figure 1(c) shows a diamond magnetic microscopy image of the same sub-region as in Fig. 1(b) under an applied field $B_{0}=30~{}{\rm mT}$. Five magnetic features are observed in the same locations as the SPION features in the SEM image. To construct a magnetic image, we use a dual-resonance optically detected magnetic resonance technique [32, 45, 46, 47], Fig. 1(d) (see also Appendix IV and Appendix V). The frequencies of two microwave tones are simultaneously discretely swept in opposite directions about both NV spin resonances. The sweeps consist of $12$ frequency steps, with a dwell time of $5.2~{}{\rm ms}$, for a total sweep length of $62~{}{\rm ms}$. The sweeps of each tone are centered about different NV spin transition frequencies, $f_{\pm}\approx D\pm\gamma_{\rm nv}B_{0}$, where $D\approx 2870~{}{\rm MHz}$ is the NV zero-field splitting and $\gamma_{\rm nv}=28.03~{}{\rm GHz/T}$ is the NV gyromagnetic ratio. Concurrently, the camera records a continuous series of frames, with exposure times synchronized to the microwave frequency steps. The directions of the microwave sweeps are periodically swapped and frames of the same microwave frequencies are averaged together to remove rolling-shutter and other timing artifacts [30]. For a given camera pixel, the intensity as a function of microwave detuning is fit to a Lorentzian function. The fitted central frequency offset, $\Delta_{c}=\gamma_{\rm nv}B_{x}$, reveals the additional magnetic field component along the NV axis due to the sample, $B_{x}$ (Appendix IV). Throughout, we define the $z$-axis to be normal to the diamond surface and the $x$-axis to be along the NV and $\vec{B_{0}}$ axes, see inset of Fig. 1(a). The fitting procedure is repeated for each pixel in the FOV in order to generate a magnetic image. Compared to sequential interrogation of the $f_{\pm}$ resonances [32, 30], this simultaneous dual-resonance method provides an increase in fluorescence contrast, with minimal change in resonance linewidth, leading to a ${\sim}1.4$-fold improvement in sensitivity.
@@
Figure 3(c) shows histograms of the fitted values of $B_{c}$ and $m_{\rm sat}$ for the $m_{x}(B_{0})$ curves of each of the 101 SPIONs. The $m_{\rm sat}$ distribution resembles a normal distribution with mean $\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$ and standard deviation $\sigma_{\rm sat}=0.12~{}{\rm A{\cdot}nm^{2}}$. The mean is consistent with, though on the lower end, of the range of values reported in the literature for similar SPIONs [51, 52, 36, 53] (Appendix VIII, Appendix VI). However, the $B_{c}$ distribution is highly asymmetric, with a peak near $B_{c}=0$ and a long tail. This reflects the presence of a large number of individual SPIONs with very sharp, step-like $m_{x}(B_{0})$ curves. The ensemble-mean behavior shown in Fig. 3(b) is more consistent with the median of this distribution, ${\rm median}(B_{c})=0.6~{}{\rm mT}$, whereas the standard deviation is much larger, $\sigma_{c}=1.4~{}{\rm mT}$.
@@
Figure 4 shows results for four example particles. SEM images of each particle are shown in Fig. 4(b). The corresponding magnetic images ($|B_{\rm hold}|=2~{}{\rm mT}$) at different times $t$ following the $B_{\rm pol}$ pulse are shown in Fig. 4(c). Each magnetic image is fit (Appendix IV.2) to extract the components of $\vec{m}(t)$ for a given particle. The quantity $\Delta m(t)\equiv|\vec{m}(t)-\vec{m}(t\rightarrow\infty)|$ is used to isolate the change in magnetic moment magnitude, $\Delta m(t)$, relative to the (often substantial) static moment, $\vec{m}(t\rightarrow\infty)$, obtained in the presence of $B_{\rm hold}$ alone, see Appendix IX. The $\Delta m(t)$ curves for each of the four example particles are shown in Fig. 4(d), along with fits to an exponential decay function. As can be seen, these four example particles span a wide range of relaxation behaviors, from nearly instantaneous ($\tau_{N}\lesssim 0.06~{}{\rm s}$) to nearly infinite ($\tau_{N}\gtrsim 20~{}{\rm s}$). Such a wide particle-to-particle variation is not necessarily surprising, given the exponential dependence of $\tau_{N}$ on SPION volume, Eq. (2), but it does validate the importance of single-particle measurements in characterizing the full sample heterogeneity.
@@ Discussion
Our results showcase the potential of diamond magnetic microscopy as a unique platform for high-throughput characterization of the magnetic properties and time-resolved dynamics of individual nanoparticles. As an early study in single-nanoparticle magnetic characterization, a number of our findings will benefit from further investigation and generalization. Examples include the sharp step-like single-SPION magnetization curves, the somewhat-low measured value of $\overline{m}_{\rm sat}$), and the faster Néel relaxation than expected from room temperature MPMS measurements.
We envision a number of improvements that can be made to the diamond magnetic microscope that may accelerate validation and future discoveries. The present magnetic image frame rate is ${\sim}16~{}{\rm Hz}$, which could be increased by several orders of magnitude using faster microwave modulation and image sensors with high-bandwidth electronics [60, 61, 62]. Combined with a reduction in coil ringdown time, either by reducing inductance or using active compensation methods [63, 64], this could enable relaxation measurements that capture $\gtrsim 6$ orders of magnitude in $\tau_{N}$ ($10^{-5}\mbox{-}10^{1}~{}{\rm s}$). The microscope’s spatial resolution and the magnetic signal strength of SPIONs can be improved by an order of magnitude each using parallel super-resolution magnetic imaging methods [36]. The magnetic sensitivity can be improved using pulsed-magnetometry protocols [65, 47] and optimized NV-doped layers [66, 67, 68]. The latter could also provide a more precise vertical distribution of NV centers, which can improve absolute quantitation, particularly if combined with the use of calibration standards [69, 70].
Future experiments may probe the correlation between SPION morphology and surface chemistry with magnetic properties, such as the anisotropy tensor, by correlating diamond magnetic microscopy measurements with high-resolution transmission electron microscopy [71, 72]. The addition of variable temperature control and higher applied magnetic fields would broaden the scope of samples that can be explored. Finally, the techniques presented here can be immediately applied to the study of a wide range of other nanomagnetic phenomena in condensed-matter and biological samples.
@@ Acknowledgements.
Author contributions. A. J., A. L., and V. M. A. conceived the idea. P. K., A. L., N. R., B. A. R., and V. M. A. designed the experiments. A. L., I. F., N. R., and B. A. R., built the main experimental apparatus. N. R. and B. A. R. acquired and analyzed magnetic microscopy and SEM data. N. R., B. A. R., and A. J. P. acquired and analyzed MPMS data. M. D. A., F. H., N. M., M. P. L., D. L. H., T. K., M. S. Z., and A. M. M. advised on SPION sample preparation and correlative electron microscopy. D. L. H. and T. K. advised on SPION magnetic properties and modeling. Y. S., J. T. D., A. J., T. K., I. F., A. B., A. L., and J. S. assisted in theoretical modeling and data interpretation. P. K., I. F., and J. S. wrote the control software. B. A. R. and N. R. wrote the initial manuscript draft. V. M. A. supervised the project. All authors helped edit the manuscript.
Funding. This work was supported by the National Science Foundation (DMR-1809800, CHE-1828744, CHE-1945148), National Institutes of Health (R21EB027405, R41MH115884, DP2GM140921), and Moore Foundation EPI-12968. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. I. F. acknowledges support from the Latvian Quantum Initiative under European Union Recovery and Resilience Facility project 2.3.1.1.i.0/1/22/I/CFLA/001.
@@ 10 $8\times 10^{12}$
Figure A6 shows the expected vertical distribution following implantation, as calculated using the Stopping Range of Ions in Matter package. To convert nitrogen to NV centers, the implanted diamonds were subsequently annealed in vacuum at $800\degree~{}{\rm C}$ for 4 hours and then $1100\degree~{}{\rm C}$ for 2 hours, following the recipe in Refs. [30, 73].
Figure A6: Simulated nitrogen depth profile. Simulated concentration of implanted nitrogen as a function of depth below the diamond surface. The simulation was performed with the Stopping and Range of Ions in Matter package.
@@ Appendix III.1 SPION sample preparation
The SPIONs used here were purchased from Ocean Nanotech (SOR-30-50). They were initially suspended in chloroform at a concentration of $25~{}{\rm mg/mL}$. Prior to depositing on a diamond, the suspension was diluted with hexane by a factor of approximately 4000 and then sonicated for 20 minutes to promote disaggregation. Then, ${\sim}0.3~{}{\rm\upmu L}$ of the diluted suspension was deposited directly on the diamond surface. The droplet’s width was larger than the diamond face’s 200 $\upmu$m dimension and smaller than its 2 mm dimension. Surface tension kept the droplet from dropping from the pipette tip from which it was dispensed, and so it was touched to the diamond face in order to transfer, and then allowed to dry. In some iterations of this sample preparation, a higher density of SPIONs was observed (using the SEM) closer to the edges of the deposited droplet, qualitatively consistent with the so-called coffee-ring phenomenon observed when nanoparticle suspensions are drop-cast [75]. The FOV containing the SPIONs studied here was located ${\sim}100~{}\upmu$m from these areas of higher SPION number density.

### v2

@@ Introduction
Here, we use diamond magnetic microscopy to image the field-dependent magnetization and time-domain Néel relaxation of over a hundred individual SPIONs in parallel. We compile histograms and extract statistics of key SPION parameters, such as saturation magnetic moment and Néel relaxation time, to characterize the ensemble distribution. The data reveal remarkable properties, including starkly steeper magnetization curves for some SPIONs compared to those obtained from ensemble methods. Such a rich heterogeneity is masked in ensemble methods, suggesting our method can produce complementary insights in fundamental studies of nanomagnetism.
@@
For diamond magnetic microscopy of SPIONs, we used the experimental setup depicted in Fig. 1(a) (see also Appendix I). The diamond is a $100\mbox{-}{\rm\upmu m}$-thick electronic-grade membrane with faces polished normal to a [110] crystal direction such that two of the four possible NV axes lie in plane. Ion implantation and annealing is used to form a ${\sim}150~{}{\rm nm}$-thick NV layer near the diamond surface on which SPION samples are dispersed (Appendix II). A beam of 532-nm light continuously illuminates NV centers in a $45\times 45~{}{\rm\upmu m}^{2}$ field of view (FOV) via an oil-immersion objective. NV center fluorescence is collected through the same objective, spectrally filtered (passing $663{\mbox{-}}800~{}{\rm nm}$), and imaged onto a scientific CMOS image sensor, with a lateral resolution of ${\sim}350~{}{\rm nm}$. The diamond is mounted on a glass coverslip using UV-curing adhesive. Microwaves are delivered to the NV centers via copper lines printed on the coverslip. The diamond is positioned so that one of the in-plane NV axes lies parallel to an applied magnetic field $\vec{B_{0}}$, generated by two pairs of coils. The diamond-coverslip assembly is mounted on a 3-axis piezo-actuated translation stage. Fluorescence images of fiducial markers milled into the diamond NV layer are used to adjust the translation stage and stabilize the diamond position [36], limiting drift to ${\lesssim}100~{}{\rm nm/day}$.
The SPIONs (Ocean Nanotech SOR30) have Fe3O4 magnetic cores and an oleic acid coating. Through transmission electron microscopy (TEM) analysis (A10), we found that the shape of the nanoparticles is predominantly spherical, with a mean diameter of $30~{}{\rm nm}$ and standard deviation of ${\sim}4~{}{\rm nm}$. Selected area electron diffraction patterns and lattice fringe spacings are consistent with the face-centered cubic Fe3O4 crystal structure. High-resolution TEM images also showed evidence that the particles are polycrystalline.
@@
The initial SPION suspensions are diluted in hexane, drop-cast onto the diamond, and dried in air, resulting in isolated single SPIONs with an average spacing of ${\sim}3~{}{\rm\upmu m}$ (Appendix III). SPIONs lying within the diamond magnetic microscope’s FOV are identified by scanning electron microscopy (SEM) and co-localized with respect to the fiducial markers. In the specific FOV presented here (Appendix III), we identified and localized $259$ single SPIONs, of which $101$ were sufficiently isolated (${\gtrsim}1.5~{}{\rm\upmu m}$ from nearest neighbor and image edges) such that their magnetic dipole patterns could be individually resolved with negligible overlap in our diamond magnetic microscope. Figure 1(b) shows a SEM image of five isolated SPIONs on the diamond surface within a sub-region of this FOV.
@@
Figure 1(c) shows a diamond magnetic microscopy image of the same sub-region as in Fig. 1(b) under an applied field $B_{0}=30~{}{\rm mT}$. Five magnetic features are observed in the same locations as the SPION features in the SEM image. To construct a magnetic image, we use a dual-resonance optically detected magnetic resonance technique [32, 45, 46, 47, 48, 49], Fig. 1(d) (see also Appendix IV and Appendix V). The frequencies of two microwave tones are simultaneously discretely swept in opposite directions about both NV spin resonances. The sweeps consist of $12$ frequency steps, with a dwell time of $5.2~{}{\rm ms}$, for a total sweep length of $62~{}{\rm ms}$. The sweeps of each tone are centered about different NV spin transition frequencies, $f_{\pm}\approx D\pm\gamma_{\rm nv}B_{0}$, where $D\approx 2870~{}{\rm MHz}$ is the NV zero-field splitting and $\gamma_{\rm nv}=28.03~{}{\rm GHz/T}$ is the NV gyromagnetic ratio. Concurrently, the camera records a continuous series of frames, with exposure times synchronized to the microwave frequency steps. The directions of the microwave sweeps are periodically swapped and frames of the same microwave frequencies are averaged together to remove rolling-shutter and other timing artifacts [30]. For a given camera pixel, the intensity as a function of microwave detuning is fit to a Lorentzian function. The fitted central frequency offset, $\Delta_{c}=\gamma_{\rm nv}B_{x}$, reveals the additional magnetic field component along the NV axis due to the sample, $B_{x}$ (Appendix IV). Throughout, we define the $z$-axis to be normal to the diamond surface and the $x$-axis to be along the NV and $\vec{B_{0}}$ axes, see inset of Fig. 1(a). The fitting procedure is repeated for each pixel in the FOV in order to generate a magnetic image. Compared to sequential interrogation of the $f_{\pm}$ resonances [32, 30], this simultaneous dual-resonance method provides an increase in fluorescence contrast, with minimal change in resonance linewidth, leading to a ${\sim}1.4$-fold improvement in sensitivity (Appendix IV.2).
@@
Figure 3(c) shows histograms of the fitted values of $B_{c}$ and $m_{\rm sat}$ for the $m_{x}(B_{0})$ curves of each of the 101 SPIONs. The $m_{\rm sat}$ distribution resembles a normal distribution with mean $\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$ and standard deviation $\sigma_{\rm sat}=0.12~{}{\rm A{\cdot}nm^{2}}$. The mean is consistent with, though on the lower end, of the range of values reported in the literature for similar SPIONs [53, 54, 36, 55] (Appendix VIII, Appendix VI). The lower $\overline{m}_{\rm sat}$ may be related to the polycrystalline structure [56] observed by high-resolution TEM (A10). Notably, the $B_{c}$ distribution is highly asymmetric, with a peak near $B_{c}=0$ and a long tail. This reflects the presence of a large number of individual SPIONs with very sharp, step-like $m_{x}(B_{0})$ curves. The ensemble-mean behavior shown in Fig. 3(b) is more consistent with the median of this distribution, ${\rm median}(B_{c})=0.6~{}{\rm mT}$, whereas the standard deviation is larger, $\sigma_{c}=1.4~{}{\rm mT}$.
@@
Figure 4 shows results for four example particles. SEM images of each particle are shown in Fig. 4(b). The corresponding magnetic images ($|B_{\rm hold}|=2~{}{\rm mT}$) at different times $t$ following the $B_{\rm pol}$ pulse are shown in Fig. 4(c). Each magnetic image is fit (Appendix IV.4) to extract the components of $\vec{m}(t)$ for a given particle. The quantity $\Delta m(t)\equiv|\vec{m}(t)-\vec{m}(t\rightarrow\infty)|$ is used to isolate the change in magnetic moment magnitude, $\Delta m(t)$, relative to the (often substantial) static moment, $\vec{m}(t\rightarrow\infty)$, obtained in the presence of $B_{\rm hold}$ alone, see Appendix IX. The $\Delta m(t)$ curves for each of the four example particles are shown in Fig. 4(d), along with fits to an exponential decay function. As can be seen, these four example particles span a wide range of relaxation behaviors, from nearly instantaneous ($\tau_{N}\lesssim 0.06~{}{\rm s}$) to nearly infinite ($\tau_{N}\gtrsim 20~{}{\rm s}$). Such a wide particle-to-particle variation is not necessarily surprising, given the exponential dependence of $\tau_{N}$ on SPION volume, Eq. (2), but it does validate the importance of single-particle measurements in characterizing the full sample heterogeneity.
@@ Discussion
Our results showcase the potential of diamond magnetic microscopy as a unique platform for high-throughput characterization of the magnetic properties and time-resolved dynamics of individual nanoparticles. As an early study in single-nanoparticle magnetic characterization, a number of our findings will benefit from further investigation and generalization. Examples include the sharp step-like single-SPION magnetization curves, the somewhat-low measured value of $\overline{m}_{\rm sat}$, and the faster Néel relaxation than expected from room temperature MPMS measurements.
We envision a number of improvements that can be made to the diamond magnetic microscope that may accelerate validation and future discoveries. The present magnetic image frame rate is ${\sim}16~{}{\rm Hz}$, which could be increased by several orders of magnitude using faster microwave modulation and image sensors with high-bandwidth electronics [63, 64, 65]. Combined with a reduction in coil ringdown time, either by reducing inductance or using active compensation methods [66, 67], this could enable relaxation measurements that capture $\gtrsim 6$ orders of magnitude in $\tau_{N}$ ($10^{-5}\mbox{-}10^{1}~{}{\rm s}$). The microscope’s spatial resolution and the magnetic signal strength of SPIONs can be improved by an order of magnitude each using parallel super-resolution magnetic imaging methods [36], though this would likely be accompanied by a slower magnetic image frame rate and lower nanoparticle characterization throughput. The magnetic sensitivity can be improved using pulsed-magnetometry protocols [68, 47] and optimized NV-doped layers [69, 70, 71]. The latter could also provide a more precise vertical distribution of NV centers, which can improve absolute quantitation, particularly if combined with the use of calibration standards [72, 73].
With advances in reproducible sample-substrate preparation [74, 75], future experiments may probe the correlation between SPION morphology and surface chemistry with magnetic properties, such as the anisotropy tensor [76], by correlating diamond magnetic microscopy measurements with high-resolution transmission electron microscopy. The addition of variable temperature control and higher applied magnetic fields would broaden the scope of samples that can be explored. Finally, the techniques presented here can be immediately applied to the study of a wide range of other nanomagnetic phenomena in condensed-matter and biological samples.
@@ Acknowledgements.
Author contributions. A. J., A. L., and V. M. A. conceived the idea. P. K., A. L., N. R., B. A. R., and V. M. A. designed the experiments. A. L., I. F., N. R., and B. A. R., built the main experimental apparatus. N. R. and B. A. R. acquired and analyzed magnetic microscopy and SEM data. N. R., B. A. R., and A. J. P. acquired and analyzed MPMS data. M. D. A., F. H., N. M., M. P. L., D. L. H., T. K., M. S. Z., and A. M. M. advised on SPION sample preparation and correlative electron microscopy. D. L. H. and T. K. advised on SPION magnetic properties and modeling. Y. S., J. T. D., A. J., T. K., I. F., A. B., A. L., and J. S. assisted in theoretical modeling and data interpretation. P. K., I. F., and J. S. wrote the control software. J. E. S. and D. E. conducted TEM imaging and analysis. B. A. R. and N. R. wrote the initial manuscript draft. V. M. A. supervised the project. All authors helped edit the manuscript.
Funding. This work was supported by the National Science Foundation (DMR-1809800, CHE-1828744, CHE-1945148), National Institutes of Health (R21EB027405, R41MH115884, DP2GM140921), and Moore Foundation EPI-12968. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. I. F. acknowledges support from the Latvian Quantum Initiative under European Union Recovery and Resilience Facility project 2.3.1.1.i.0/1/22/I/CFLA/001. The research at University of Nebraska-Lincoln was performed in part in the Nebraska Nanoscale Facility: National Nanotechnology Coordinated Infrastructure by the National Science Foundation under award no. ECCS: 2025298, and with support from the Nebraska Research Initiative through the Nebraska Center for Materials and Nanoscience.
@@
Only one NV axis (out of four possible axes) is used for magnetic measurements. We used a [110]-cut diamond, where two of the NV axes are in plane (i.e. the plane of the largest diamond faces). We selected one of the these in-plane NV axes for magnetic measurements. The diamond was positioned so that this axis was parallel to the applied magnetic field. The laser beam was linearly polarized. A half waveplate was used to align the polarization axis approximately perpendicular to the selected NV axis to maximize ODMR contrast.
@@ 10 $8\times 10^{12}$
Figure A6 shows the expected vertical distribution following implantation, as calculated using the Stopping Range of Ions in Matter (SRIM) package. To convert nitrogen to NV centers, the implanted diamonds were subsequently annealed in vacuum at $800\degree~{}{\rm C}$ for 4 hours and then $1100\degree~{}{\rm C}$ for 2 hours, following the recipe in Refs. [30, 77].
Figure A6: Simulated nitrogen depth profile. The concentration of implanted nitrogen is plotted as a function of depth below the diamond surface. The simulation was performed with the SRIM package.
@@ Appendix III.1 SPION sample preparation
The SPIONs used here were purchased from Ocean Nanotech (SOR30). They were initially suspended in chloroform at a concentration of $25~{}{\rm mg/mL}$. Prior to depositing on a diamond, the suspension was diluted with hexane by a factor of approximately 4000 and then sonicated for 20 minutes to promote disaggregation. Then, ${\sim}0.3~{}{\rm\upmu L}$ of the diluted suspension was deposited directly on the diamond surface. The droplet’s diameter was larger than the diamond face’s $200~{}{\rm\upmu m}$ dimension and smaller than its $2~{}{\rm mm}$ dimension. Surface tension kept the droplet from dropping from the pipette tip from which it was dispensed, and so it was touched to the diamond face in order to transfer, and then allowed to dry. In some iterations of this sample preparation, a higher density of SPIONs was observed (using the SEM) closer to the edges of the deposited droplet, qualitatively consistent with the so-called coffee-ring phenomenon observed when nanoparticle suspensions are drop-cast [79]. The FOV containing the SPIONs studied here was located ${\sim}100~{}\upmu$m from these areas of higher SPION number density.